OpenStudy (anonymous):

Let x and y be two numbers. Find the solution of the puzzle. (in the comments)

OpenStudy (anonymous):



OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

The difference of 3y and 12x is 1. 3y - 2x = 1

OpenStudy (saifoo.khan):

y is 1/3 more than 4 times the value of x. I'm not sure about the first one. It can be: y + 1/3 = 4x

OpenStudy (anonymous):

So.. what do I do with it?

OpenStudy (saifoo.khan):

@yummydum : Apply your brilliant brain over here.

OpenStudy (saifoo.khan):

C'mon. I didn't asked you to write a novel. -_-

OpenStudy (anonymous):

\[y=4x+{1\over3}\]\[3x-12y=1\]substitute the value of y:\[3x-12(4x+{1\over3})=1\]distribute the -12:\[3x-48x-4=1\]add 4 to both sides:\[3x-48x=5\]combine the like terms:\[-45x=5\]divide both sides by -45:\[x=-9\] now plug that value into the first equation:\[y=4(-9)+{1\over3}\]distribute:\[y=-36+{1\over3}\]add:\[y=36{1\over3}\] if those dont work try flipping the 3x and 12x in the second equation :)

OpenStudy (saifoo.khan):

Nice use of your brown brain. @yummydum

OpenStudy (anonymous):

bakol tumhare mere paas koi brain naam ki cheez hi nahi :P

OpenStudy (saifoo.khan):

Usme koi shaq nahi hai. ^_~ But thora bohat Pakistani honay ki wajah say izat tou rakhni parti haina. :P

OpenStudy (anonymous):

rakho rakho :P

OpenStudy (saifoo.khan):

Sure sure.

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

Your Welcome! :D